Circular No. 8042

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003B NEAR NGC 1097 R. Evans, Hazelbrook, N.S.W., reports his visual discovery of an ap-

parent supernova (mag ∼ 15.0) on Jan. 5.5 UT with a 0.41-m reflector at Frogrock, near Mudgee, N.S.W. Evans estimates the new object to be 90′′ west and 175′′ north of the nucleus of NGC 1097 (α = 2^{h}46^{m}.3, δ = −30^{o}17′, equinox 2000.0) in a faint emission region visible on the European Southern Observatory B Survey image (and he adds that SN 2003B is also located 40′′ due east of the nearby small elliptical galaxy NGC 1097A); no star is seen at this location on either the ESO B or J Survey. The new object was confirmed visually by S. Quirk.

R. Kirshner and J. Silverman, Harvard-Smithsonian Center for Astrophysics, obtained a spectrogram (range 400–900 nm) of SN 2003B on Jan. 6.17 UT with the Clay 6.5-m telescope at Las Campanas Observatory. The spectrum resembles that of a type-II supernova, < 2 weeks past maximum. A well-developed P-Cyg line at Hα has a peak near the systemic velocity of NGC 1097, which is marked by a narrow emission, presumably due to the H II region near the event. The deepest absorption of the P-Cyg line at Hα is shifted ∼ 6300 km/s toward the blue.

A. V. Filippenko and R. Chornock, University of California, Berkeley, write: “Inspection of CCD spectra (range 310–1000 nm) obtained on Jan. 6 UT with the Keck I 10-m telescope (+ LRIS) shows that SN 2003B is of type II, a few weeks past the explosion. Hα exhibits a well-developed P-Cyg profile, with an absorption minimum corresponding to an expansion velocity of 7000 km/s (assuming the NED redshift of NGC 1097, 1275 km/s). The expansion velocity measured from Hβ is 5600 km/s. Many Fe II and other absorption lines are also visible. From our data, it is not possible to determine whether the supernova is in NGC 1097 or 1097A, whose redshifts are nearly identical.”

SUPERNOVA 2003A IN UGC 5904 Filippenko and Chornock also report that inspection of CCD spectra,

obtained as above on Jan. 6 UT, shows that SN 2003A (IAUC 8041) is of type Ib/c, probably a few weeks after maximum brightness. Weak He I absorption lines are visible, but the overall spectrum resembles that of type-Ic supernovae. The redshift derived from a superposed H II region is ∼ 7000 km/s. The supernova is probably a member of the southern spiral galaxy in the interacting pair UGC 5904, even though it appears projected closer to the nucleus of the northern elliptical galaxy.

2003 January 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green